$SPAD/src/input perman.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool perman.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
kn n ==
  r : MATRIX INT := new(n,n,1)
  for i in 1..n repeat
    r.i.i := 0
  r
--R
--R                                                                   Type: Void
--E 1
--S 2 of 3
permanent(kn(5) :: SQMATRIX(5,INT))
--R
--R   Compiling function kn with type PositiveInteger -> Matrix(Integer)
--R
--R   (2)  44
--R                                                        Type: PositiveInteger
--E 2
--S 3 of 3
[permanent(kn(n) :: SQMATRIX(n,INT)) for n in 1..13]
--R
--R   Cannot compile conversion for types involving local variables. In
--R      particular, could not compile the expression involving ::
--R      SQMATRIX(n,INT)
--R   AXIOM will attempt to step through and interpret the code.
--R
--R   (3)
--R   [0,1,2,9,44,265,1854,14833,133496,1334961,14684570,176214841,2290792932]
--R                                               Type: List(NonNegativeInteger)
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4